OpenStudy (anonymous):

f(x)=-2x^3+5, find x if f(x)=-49

OpenStudy (asevilla5):

are there any parentheses in the equation?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

you need to solve \(-2x^3+5=-49\)

OpenStudy (zzr0ck3r):

ubtract 5 from both sides divide both sides by -2 cube root both sides in that order

OpenStudy (anonymous):

o equate both of the given functions...as -2x^3+5=-49 =>-2(x^3)+5-5=-49-5 =>-2x^3=-54 =>x^3=54/2=27 =>x= cuberoot 27= (3^3)^1/3=3^(3/3)=3 so x=3

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you are welcome.